# Exactly solvable dissipative dynamics and one-form strong-to-weak spontaneous symmetry breaking in interacting two-dimensional spin systems

### v1

@@ Abstract
We study the dissipative dynamics of a class of interacting “gamma-matrix” spin models coupled to a Markovian environment. For spins on an arbitrary graph, we construct a Lindbladian that maps to a non-Hermitian model of free Majorana fermions hopping on the graph with a background classical $\mathbb{Z}_{2}$ gauge field. We show, analytically and numerically, that the steady states and relaxation dynamics are qualitatively independent of the choice of the underlying graph, in stark contrast to the Hamiltonian case. We also show that the exponentially many steady states provide a concrete example of mixed-state topological order, in the sense of strong-to-weak spontaneous symmetry breaking of a one-form symmetry. While encoding only classical information, the steady states still exhibit long-range quantum correlations. Afterward, we examine the relaxation processes toward the steady state by numerically computing decay rates, which we generically find to be finite, even in the dissipationless limit. We however identify symmetry sectors where fermion-parity conservation is enhanced to fermion-number conservation, where we can analytically bound the decay rates and prove that they vanish in the limits of both infinitely weak and infinitely strong dissipation. Finally, we show that while the choice of coherent dynamics is very flexible, exact solvability strongly constrains the allowed form of dissipation. Our work establishes an analytically tractable framework to explore nonequilibrium quantum phases of matter and the relaxation mechanisms toward them.
@@ I Introduction
Interacting spin systems in two dimensions (2D) are a fertile ground for fundamental and applied physics. Among the many phases of matter that 2D spin systems can host, quantum spin liquids (QSL) [1, 2] stand out due to their exotic properties, such as the absence of local magnetic ordering, fractionalized excitations, and often topological order [3, 4, 5, 6, 7, 8]. These properties also make QSLs and other topologically ordered states potential candidates to realize fault-tolerant quantum computers and topological quantum memories [9, 10, 11]. These fundamental and potentially technological features motivate intense theoretical and experimental research into QSLs [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
@@
In the rich landscape of QSLs, the Kitaev honeycomb model [8]—a model of spins-$1/2$ on the honeycomb lattice with highly anisotropic Heisenberg interactions—occupies a special place, as it provides a rare example of an exactly solvable model with both gapped and gapless QSL phases. The fractionalization of the spins into free itinerant Majorana fermions and static classical $\mathbb{Z}_{2}$ gauge fields yields not only a detailed knowledge of the ground states but also key insights into the dynamics and signatures of the anyonic excitations [23, 24, 25, 26, 27, 28, 29, 21, 22]. Several variants of the Kitaev model that preserve exact solvability while hosting distinct types of QSLs have been proposed [30, 31, 32, 33, 34]. However, imperfections in devices or samples and external influences are, to a certain extent, unavoidable in realistic systems, including in current-term noisy quantum hardware. It is thus natural to ask what the fate of topological order and the QSL state is once the system is explicitly coupled to the environment: which, if any, exactly solvable QSL models can still be constructed and what is their phenomenology?
@@
In this paper, we show that Kitaev’s approach can be used to construct an exactly solvable interacting model on an arbitrary two-dimensional graph, while including coupling to a Markovian bath that leads to Lindbladian dynamics. The local degrees of freedom of these models are “gamma matrices” [33, 31, 32], replacing Pauli matrices of spin-1/2 models by elements of a suitable Clifford algebra. In stark contrast to the closed-system limit, where qualitatively different states arise for different choices of 2D spin models [8, 30, 31, 32], we show that the qualitative properties of these models are largely insensitive to details of the coherent interactions (e.g., graph geometry or relative coupling strengths). In particular, the steady-state manifold is completely determined by symmetry and has the same structure for any graph or parameters of the model. The transient dynamics are also qualitatively similar. Two special cases of our model on the square lattice were studied in Refs. 35, 36, 37. There, the authors identified the steady states, reduced the Lindbladian to its free-fermion representation inside fixed symmetry sectors, and numerically studied the dominant relaxation timescales. Other graphs have not been investigated; most prominently missing is a free-fermion dissipative honeycomb model (the spin-$1/2$ honeycomb model of Ref. 38, which, while sharing the same steady states, is not free-fermion reducible).
@@
In addition to providing key insights into the dynamics of dissipative QSLs, the gamma-matrix model is also particularly useful to study the intensely debated nature of mixed-state topological order [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. In the ground-state physics of pure states, the characteristic features of topological order, such as robust ground-state degeneracy and long-range correlations, come hand-in-hand with the ability to store and manipulate quantum information coherently. On the other hand, it has been understood that mixed states are richer and allow exotic phases of matter where a quantum memory is degraded to a classical memory but still exhibits nontrivial long-range quantum correlations [52, 53, 50, 51, 54]. It was also recognized [55] that the transition between a coherent ground-state phase and an incoherent mixed-state phase can happen sharply at a critical decoherence strength (or, in a continuous-time description, dynamically at a critical time), when a strong symmetry of the system is spontaneously broken [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 54, 70, 71, 72, 73, 74, 75] to a weak symmetry. If the spontaneously broken symmetry is a higher-form symmetry [59], i.e., whose symmetry operators and charged operators are extended objects, then the resulting mixed state exhibits a sense of mixed-state topological order. We show that the gamma-matrix models give examples of such strong-to-weak spontaneous symmetry breaking (SW-SSB) and that its steady-state manifold forms a classical memory with such quantum mixed-state topological order.
@@
Figure 1: Schematic representation of the main results. (a) A “proper edge coloring” of a graph (i.e., such that no two edges of the same color meet at any vertex) defines a gamma-matrix Hamiltonian with Heisenberg interactions set by the edge colors. The gamma-matrix “spins” and itinerant Majorana fermions (resulting from the gamma-matrix-to-Majorana mapping of Sec. IV) live on the vertices and gauge fields $u_{ij}$ (resulting from the same mapping) live on the edges. The gauge fields give rise to fluxes $W_{p}$ through elementary plaquettes $p$. (b) When the system is realized on a topologically nontrivial surface like a torus, fluxes can thread noncontractible loops. One can define logical operators $\overline{X}_{a}$ and $\overline{Z}_{a}$ for each noncontractible loop. In the steady state, the product of two $\overline{Z}$ operators acquires a nonzero correlation even when the two operators are arbitrarily far apart. Consequently, the $\overline{X}$ operators are spontaneously broken from a strong to a weak one-form symmetry. (c) Vectorization: doubling of degrees of freedom to represent the open system. In the free-fermion solvable case, two copies (layers) of the graph are connected by interlayer gauge fields $v_{j}$ (black vertical lines) arising from dephasing. The symmetry sectors of the Lindbladian are labeled by $W_{p}$, $\overline{X}_{a}$, and interlayer fluxes $\mathcal{W}_{q}$, and each has its own fermionic vacuum. The steady states are the fermionic vacua in sectors with $\mathcal{W}_{q}=1$ for all interlayer plaquettes $q$, i.e., without any gauge flips ($\mathcal{W}_{q}=-1)$, and are labeled by their flux patterns. The decaying states of the Lindbladian are the vacua with gauge flips and fermionic excitations. (d) When intralayer gauge fields are flipped, the excitations conserve only fermionic parity, which is determined by a Pfaffian method we describe. For weak dissipation, their decay rate is proportional to the Hamiltonian coupling strength $J$ when the thermodynamic limit is taken first (anomalous relaxation). (d) If only interlayer gauge fields $v_{j}$ are flipped, the excitations are labeled by a conserved fermion number and their decay rates can be analytically bounded with Bendixson inequalities (shaded region). At weak dissipation, the decay rate of all excitations is linear in the dissipation strength $\gamma$.
@@ II.1 The Kitaev honeycomb model
The Pauli matrices describing spin-$1/2$ degrees of freedom are the simplest (two-dimensional) Clifford algebra: the two generators $\sigma^{x}_{i}=(\sigma^{x}_{i})^{\dagger}$ and $\sigma^{y}_{i}=(\sigma^{y}_{i})^{\dagger}$ satisfy $\{\sigma^{x}_{i},\sigma^{y}_{i}\}=0$ and $(\sigma^{x}_{i})^{2}=(\sigma^{y}_{i})^{2}=1$, and $\sigma^{z}_{i}=-\mathrm{i}\sigma^{x}_{i}\sigma^{y}_{i}$ is the chiral element of the algebra (i.e., an operator that anticommutes with all its elements). The crucial feature of Eq. (1) is that there is a single $x$-, $y$-, and $z$-type link emanating from each site, which allowed Kitaev to reduce the Hamiltonian to that of free Majorana fermions hopping on a background gauge field [8], after fractionalization of the spins into Majorana fermions. For lattices with a higher valence (site coordination number), the same construction is still possible if one chooses the local degrees of freedom as gamma matrices in a higher-dimensional Clifford algebra. Hamiltonians of this type were studied in Refs. 33, 31, 32.
@@ II.2 The gamma-matrix model
We consider a model defined on an arbitrary graph $G$ with $N$ vertices $i$ and $E$ edges $\langle ij\rangle$, with $z=z(i)$ the local vertex valence (number of incident edges at vertex $i$). The local degrees of freedom are Euclidean gamma matrices $\Gamma_{j}^{\mu}=(\Gamma_{j}^{\mu})^{\dagger}$, $j\in\{1,\dots,N\}$ and $\mu\in\{1,\dots,2k\}$, of dimension $2^{k}\times 2^{k}$ that satisfy $\{\Gamma_{j}^{\mu},\Gamma_{j}^{\nu}\}=2\delta_{\mu\nu}$ and $[\Gamma^{\mu}_{i},\Gamma^{\nu}_{j}]=0$ for $i\neq j$. Here, $2k=z$ ($12k=z+$ 1) if $z$ is even (odd). Moreover, for convenience, we denote $\Gamma_{j}^{0}=\mathbbm{1}$ (noting that $\Gamma^{0}_{j}$ is not a member of the Clifford algebra and does not satisfy the anticommutation relations). The set of all possible products of gamma matrices forms a basis of operators of the local Hilbert space; of particular importance are the $\mu\neq\nu$ bilinears,
@@
Figure 2: Spin Hamiltonians defined by the color rule. Each edge of the graph is assigned a color and two gamma matrices with that color at the two endpoints of the edge interact in a term of the Hamiltonian. We illustrate the general construction for a vertex $i$ of valence $z(i)=5$.
@@ II.3 Symmetries of the gamma-matrix Hamiltonian
Not all strings of the form (6) commute with the Hamiltonian. Since each $K_{ij}$ in $H$ anticommutes with precisely zero or two $K_{i^{\prime}j^{\prime}}$ in $W_{\mathcal{C}}$, we find $[H,W_{\mathcal{C}}]=0$ for all closed $\mathcal{C}$. On the other hand, the $K_{ij}$ at the endpoint of an open string $W_{\mathcal{P}}$ anticommutes with all $K_{i^{\prime}j^{\prime}}$ that share that endpoint but have a different color; hence, $[H,W_{\mathcal{P}}]\neq 0$ for all open $\mathcal{P}$. The addition of a gamma matrix $\Gamma_{j}^{\mu}$ with $\mu\in\{1,\dots,2k\}$ at the end of the string leads to commutation with all $K_{i^{\prime}j^{\prime}}$ except the one with color $\mu$; hence, also $[H,W_{\mathcal{P}}^{\mu\nu}]\neq 0$ for all open $\mathcal{P}$ and $\mu,\nu\in\{1,\dots,2k\}$. Finally, open string operators commute with all $K_{ij}$, and hence the Hamiltonian, if and only if they terminate at a chiral element at both ends; hence, $[H,W_{\mathcal{P}}^{\mathrm{ch}}]=0$.
@@ II.4 Lindbladian dynamics
Our main focus will be on free-fermion reducible Lindbladians (see Sec. IV), which, as we shall show, requires single-site generalized “dephasing” jump operators (see App. A for further details):
@@
where $\gamma_{j}$ is the on-site dissipation strength. [^1] If $z(j)$ is odd, $\Gamma_{j}^{2k}$ enters neither the Hamiltonian nor the jump operators. We can either regard it as an inert gamma matrix only necessary to complete the Hilbert space—which replicates the Lindbladian, leading to two degenerate copies that can be treated individually; or add a jump operator $L^{\prime}_{j}=\sqrt{\gamma^{\prime}_{j}}\Gamma^{2k}_{j}$—the formulae below must then be modified accordingly but the qualitative picture remains intact. (We use the terms dissipation and dephasing interchangeably in this work.) More generally, the symmetry analysis of Secs. II.6 and III holds for any jump operator that is a single string of gamma matrices and conserves all fluxes $W_{\mathcal{C}}$, i.e., $[L_{m},W_{\mathcal{C}}]=0$ for all $m$ and $\mathcal{C}$. An example of such a jump operator is the edge operator $L_{\langle ij\rangle}\propto K_{ij}$ [38]. This choice does not, however, preserve the free-fermion description of the model (see App. A). Simultaneous flux conservation and free-fermion solvability reduce the choice of jump operators to Eq. (11).

### v2

@@ Abstract
We study the dissipative dynamics of a class of interacting “gamma-matrix” spin models coupled to a Markovian environment. For spins on an arbitrary graph, we construct a Lindbladian that maps to a non-Hermitian model of free Majorana fermions hopping on the graph with a background classical $\mathbb{Z}_{2}$ gauge field. We show, analytically and numerically, that the steady states and relaxation dynamics are qualitatively independent of the choice of the underlying graph, in stark contrast to the Hamiltonian case. We also show that the exponentially many steady states provide a concrete example of mixed-state topological order, in the sense of strong-to-weak spontaneous symmetry breaking of a one-form symmetry. While encoding only classical information, the steady states still exhibit long-range quantum correlations. Afterward, we examine the relaxation processes toward the steady state by numerically computing decay rates, which we generically find to be finite, even in the dissipationless limit. However, we identify symmetry sectors where fermion-parity conservation is enhanced to fermion-number conservation, where we can analytically bound the decay rates and prove that they vanish in the limits of both infinitely weak and infinitely strong dissipation. Finally, we show that while the choice of coherent dynamics is very flexible, exact solvability strongly constrains the allowed form of dissipation. Our work establishes an analytically tractable framework to explore nonequilibrium quantum phases of matter and the relaxation mechanisms toward them.
@@ I Introduction
Interacting spin systems in two dimensions (2D) are a fertile ground for fundamental and applied physics. Among the many phases of matter that 2D spin systems can host, quantum spin liquids (QSLs) [1, 2] stand out due to their exotic properties, such as the absence of local magnetic ordering, fractionalized excitations, and often topological order [3, 4, 5, 6, 7, 8]. These properties also make QSLs and other topologically ordered states potential candidates to realize fault-tolerant quantum computers and topological quantum memories [9, 10, 11]. These fundamental and potentially technological features motivate intense theoretical and experimental research into QSLs [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
@@
In the rich landscape of QSLs, the Kitaev honeycomb model [8]—a model of spins-$1/2$ on the honeycomb lattice with highly anisotropic Heisenberg interactions—occupies a special place, as it provides a rare example of an exactly solvable model with both gapped and gapless QSL phases. The fractionalization of the spins into free itinerant Majorana fermions and static classical $\mathbb{Z}_{2}$ gauge fields yields not only a detailed knowledge of the ground states but also key insights into the dynamics and signatures of the anyonic excitations [23, 24, 25, 26, 27, 28, 29, 21, 22]. Several variants of the Kitaev model that preserve exact solvability while hosting distinct types of QSLs have been proposed [30, 31, 32, 33, 34]. However, imperfections in devices or samples and external influences are, to a certain extent, unavoidable in realistic systems, including in current-term noisy quantum hardware. It is thus natural to ask what the fate of topological order and the QSL state is once the system is explicitly coupled to the environment: which, if any, exactly solvable QSL models can still be constructed, and what is their phenomenology?
@@
In this paper, we show that Kitaev’s approach can be used to construct an exactly solvable interacting model on an arbitrary two-dimensional graph, while including coupling to a Markovian bath that leads to Lindbladian dynamics. The local degrees of freedom of these models are “gamma matrices” [33, 31, 32], replacing Pauli matrices of spin-1/2 models by elements of a suitable Clifford algebra. In stark contrast to the closed-system limit, where qualitatively different states arise for different choices of 2D spin models [8, 30, 31, 32], we show that the qualitative properties of these models are largely insensitive to details of the coherent interactions (e.g., graph geometry or relative coupling strengths). In particular, the steady-state manifold is completely determined by symmetry and has the same structure for any graph or parameters of the model. The transient dynamics are also qualitatively similar. Two special cases of our model on the square lattice were studied in Refs. 35, 36, 37. There, the authors identified the steady states, reduced the Lindbladian to its free-fermion representation inside fixed symmetry sectors, and numerically studied the dominant relaxation timescales. Other graphs have not been investigated; most prominently missing is a free-fermion dissipative honeycomb model (the spin-$1/2$ honeycomb model of Ref. 38, which, while sharing the same steady states, is not free-fermion-reducible; see also Ref. 39).
@@
In addition to providing key insights into the dynamics of dissipative QSLs, the gamma-matrix model is also particularly useful to study the intensely debated nature of mixed-state topological order [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. In the ground-state physics of pure states, the characteristic features of topological order, such as robust ground-state degeneracy and long-range correlations, go hand-in-hand with the ability to store and manipulate quantum information coherently. On the other hand, it has been understood that mixed states are richer and allow exotic phases of matter where a quantum memory is degraded to a classical memory but still exhibits nontrivial long-range quantum correlations [53, 54, 51, 52, 55]. It was also recognized [56] that the transition between a coherent ground-state phase and an incoherent mixed-state phase can happen sharply at a critical decoherence strength (or, in a continuous-time description, dynamically at a critical time), when a strong symmetry of the system is spontaneously broken [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 55, 71, 72, 73, 74, 75, 76] to a weak symmetry. If the spontaneously broken symmetry is a higher-form symmetry [60], i.e., whose symmetry operators and charged operators are extended objects, then the resulting mixed state exhibits a sense of mixed-state topological order. We show that the gamma-matrix models give examples of such strong-to-weak spontaneous symmetry breaking (SW-SSB) and that its steady-state manifold forms a classical memory with such quantum mixed-state topological order.
@@
Figure 1: Schematic representation of the main results. (a) A “proper edge coloring” of a graph (i.e., such that no two edges of the same color meet at any vertex) defines a gamma-matrix Hamiltonian with Heisenberg interactions set by the edge colors. The gamma-matrix “spins” and itinerant Majorana fermions (resulting from the gamma-matrix-to-Majorana mapping of Sec. IV) live on the vertices and gauge fields $u_{ij}$ (resulting from the same mapping) live on the edges. The gauge fields give rise to fluxes $W_{p}$ through elementary plaquettes pp. (b) When the system is realized on a topologically nontrivial surface like a torus, fluxes can thread noncontractible loops. One can define logical operators $\overline{X}_{a}$ and $\overline{Z}_{a}$ for each noncontractible loop. In the steady state, the product of two $\overline{Z}$ operators acquires a nonzero correlation even when the two operators are arbitrarily far apart. Consequently, the $\overline{X}$ operators are spontaneously broken from a strong to a weak one-form symmetry. (c) Vectorization: doubling of degrees of freedom to represent the open system. In the free-fermion solvable case, two copies (layers) of the graph are connected by interlayer gauge fields $v_{j}$ (black vertical lines) arising from dephasing. The symmetry sectors of the Lindbladian are labeled by $W_{p}$, $\overline{X}_{a}$, and interlayer fluxes $\mathcal{W}_{q}$, and each has its own fermionic vacuum. The steady states are the fermionic vacua in sectors with $\mathcal{W}_{q}=1$ for all interlayer plaquettes $q$, i.e., without any gauge flips ($\mathcal{W}_{q}=-1)$, and they are labeled by their flux patterns. The decaying states of the Lindbladian are the vacua with gauge flips and fermionic excitations. (d) When intralayer gauge fields are flipped, the excitations conserve only fermionic parity, which is determined by a Pfaffian method we describe. For weak dissipation, their decay rate is proportional to the Hamiltonian coupling strength $J$ when the thermodynamic limit is taken first (anomalous relaxation). (d) If only interlayer gauge fields $v_{j}$ are flipped, the excitations are labeled by a conserved fermion number, and their decay rates can be analytically bounded with Bendixson inequalities (shaded region). At weak dissipation, the decay rate of all excitations is linear in the dissipation strength $\gamma$.
@@ II.1 The Kitaev honeycomb model
The Pauli matrices describing spin-$1/2$ degrees of freedom are the simplest (two-dimensional) Clifford algebra: the two generators $\sigma^{x}_{i}=(\sigma^{x}_{i})^{\dagger}$ and $\sigma^{y}_{i}=(\sigma^{y}_{i})^{\dagger}$ satisfy $\{\sigma^{x}_{i},\sigma^{y}_{i}\}=0$ and $(\sigma^{x}_{i})^{2}=(\sigma^{y}_{i})^{2}=1$, and $\sigma^{z}_{i}=-\mathrm{i}\sigma^{x}_{i}\sigma^{y}_{i}$ is the chiral element of the algebra (i.e., an operator that anticommutes with all its elements). The crucial feature of Eq. (1) is that there is a single $x$-, $y$-, and $z$-type link emanating from each site, which allowed Kitaev to reduce the Hamiltonian to that of free Majorana fermions hopping on a background gauge field [8], after fractionalization of the spins into Majorana fermions. (Some variants of the Kitaev model can also be solved without recourse to fractionalization [125, 126].) For lattices with a higher valence (site coordination number), the same construction is still possible if one chooses the local degrees of freedom as gamma matrices in a higher-dimensional Clifford algebra. Hamiltonians of this type were studied in Refs. 33, 31, 32.
@@ II.2 The gamma-matrix model
We consider a model defined on an arbitrary graph $G$ with $N$ vertices $i$ and $E$ edges $\langle ij\rangle$, with $z=z(i)$ the local vertex valence (number of incident edges at vertex $i$). The local degrees of freedom are Euclidean gamma matrices $\Gamma_{j}^{\mu}=(\Gamma_{j}^{\mu})^{\dagger}$, $j\in\{1,\dots,N\}$ and $\mu\in\{1,\dots,2k\}$, of dimension $2^{k}\times 2^{k}$ that satisfy $\{\Gamma_{j}^{\mu},\Gamma_{j}^{\nu}\}=2\delta_{\mu\nu}$ and $\Gamma^{\mu}_{i},\Gamma^{\nu}_{j}]=0$ for $i\neq j$. Here, $2k=z$ ($2k=z+1$) if $z$ is even (odd). Moreover, for convenience, we denote $\Gamma_{j}^{0}=\mathbbm{1}$ (noting that $\Gamma^{0}_{j}$ is not a member of the Clifford algebra and does not satisfy the anticommutation relations). The set of all possible products of gamma matrices forms a basis of operators of the local Hilbert space; of particular importance are the $\mu\neq\nu$ bilinears,
@@
Figure 2: Spin Hamiltonians defined by the color rule. Each edge of the graph is assigned a color, and two gamma matrices with that color at the two endpoints of the edge interact in a term of the Hamiltonian. We illustrate the general construction for a vertex $i$ of valence $z(i)=5$.
@@ II.3 Symmetries of the gamma-matrix Hamiltonian
Not all strings of the form (6) commute with the Hamiltonian. Since each $K_{ij}$ in $H$ anticommutes with precisely zero or two $K_{i^{\prime}j^{\prime}}$ in $W_{\mathcal{C}}$, we find $H,W_{\mathcal{C}}]=0$ for all closed $\mathcal{C}$. On the other hand, the $K_{ij}$ at the endpoint of an open string $W_{\mathcal{P}}$ anticommutes with all $K_{i^{\prime}j^{\prime}}$ that share that endpoint but have a different color; hence, $H,W_{\mathcal{P}}]\neq 0$ for all open $\mathcal{P}$. The addition of a gamma matrix $\Gamma_{j}^{\mu}$ with $\mu\in\{1,\dots,2k\}$ at the end of the string leads to commutation with all $K_{i^{\prime}j^{\prime}}$ except the one with color $\mu$; hence, also $H,W_{\mathcal{P}}^{\mu\nu}]\neq 0$ for all open $\mathcal{P}$ and $\mu,\nu\in\{1,\dots,2k\}$. Finally, open string operators commute with all $K_{ij}$, and hence the Hamiltonian, if and only if they terminate at a chiral element at both ends; hence, $H,W_{\mathcal{P}}^{\mathrm{ch}}]=0$.
@@ II.4 Lindbladian dynamics
Our main focus will be on free-fermion reducible Lindbladians (see Sec. IV), which, as we shall show, requires single-site generalized “dephasing” jump operators (see Appendix A for further details):
@@
where $\gamma_{j}$ is the on-site dissipation strength. [^1] If $z(j)$ is odd, $\Gamma_{j}^{2k}$ enters neither the Hamiltonian nor the jump operators. We can either regard it as an inert gamma matrix only necessary to complete the Hilbert space—which replicates the Lindbladian, leading to two degenerate copies that can be treated individually; or add a jump operator $L^{\prime}_{j}=\sqrt{\gamma^{\prime}_{j}}\Gamma^{2k}_{j}$—the formulas below must then be modified accordingly but the qualitative picture remains intact. (We use the terms dissipation and dephasing interchangeably in this work.) The jump operators (11) generalize the standard dephasing mechanism in spin-$1/2$ systems ($L_{j}=\sqrt{\gamma_{j}}\sigma^{z}_{j}$) to gamma-matrix “spins”. More generally, the symmetry analysis of Secs. II.6 and III holds for any jump operator that is a single string of gamma matrices and conserves all fluxes $W_{\mathcal{C}}$, i.e., $L_{m},W_{\mathcal{C}}]=0$ for all $m$ and $\mathcal{C}$. An example of such a jump operator is the edge operator $L_{\langle ij\rangle}\propto K_{ij}$ [38]. This choice does not, however, preserve the free-fermion description of the model (see Appendix A). Simultaneous flux conservation and free-fermion solvability reduce the choice of jump operators to Eq. (11). We will show below that, once this choice is made, the dynamics of the open system are largely independent of the choice of graph (i.e., of Hamiltonian); in contrast, for closed systems, different choices of graph lead to vastly different physics [8, 30, 31, 32, 33, 34].